# LT-Soups: Bridging Head and Tail Classes via Subsampled Model Soups

### v1

@@ 5 Experiments
We evaluate our method on both synthetically constructed and naturally occurring long-tailed (LT) datasets. For synthetic benchmarks, we use CIFAR-100-LT, ImageNet-LT, and Places-LT—long-tailed variants derived from their balanced counterparts by sampling class instances according to Pareto or exponential decay distributions liu2019large. These datasets exhibit sample counts ranging from 1,280 to as few as 5 images per class. For real-world evaluation, we include iNaturalist 2018 (8,142 classes, 437.5K images) and NIH-CXR-LT (20 classes, 88.5K images), which reflect different imbalance structures, with approximately 10% and 90% head classes, respectively. To assess performance across the long-tail spectrum, we also report the average accuracy across all five datasets. Following liu2019large, we evaluate separately on many-shot (>>100 samples), medium-shot (20–100), and few-shot (<<20) class subsets. For ablation analysis, we use TinyImageNet-LT, which contains 200 classes with sample counts ranging from 500 in head classes to 5 in tail classes. To conserve space, we present only CLIP-based results in the main text; additional implementation details and extended results are included in Appendix D.
@@
Methods CIFAR100-LT Places-LT ImageNet-LT $\rho$=100 $\eta$=0.54 $\rho$=996 $\eta$=0.55 $\rho$=256 $\eta$=0.62 All Many Med. Few All Many Med. Few All Many Med. Few BALLAD ma2021simple - - - - 49.5 49.3 50.2 48.4 75.7 79.1 74.5 69.8 Decoder wang2024exploring - - - - 46.8 - - - 73.2 - - - LPT dong2022lpt - - - - 50.1 49.3 52.3 46.9 - - - - Linear Probing 70.0 77.2 71.1 60.4 48.8 48.8 49.7 47.1 74.2 77.8 73.3 67.4 Full-FT 79.6 88.1 79.9 69.3 46.6 49.9 46.3 41.4 73.9 79.8 71.9 63.9 cRT kang2019decoupling 78.8 89.7 79.7 65.1 44.4 51.0 43.1 35.4 72.6 81.1 70.6 56.1 PEFT shi2024long 81.3 85.2 80.9 77.1 51.5 51.3 52.2 50.5 77.0 80.2 76.1 71.5 Model Soups wortsman2022model 82.1 89.9 82.2 73.0 49.4 51.7 50.0 43.7 76.0 81.5 74.5 65.5 LT-Soups (Ours) 83.5 88.2 83.5 78.0 51.7 51.2 52.8 50.3 77.4 81.2 76.1 70.7
@@
Methods NIH-CXR-LT iNaturalist 2018 $\rho$=6491 $\eta$=5.66 $\rho$=500 $\eta$=0.11 All Many Med. Few All Many Med. Few BALLAD ma2021simple 34.5 36.7 38.9 20.8 49.5 49.3 50.2 48.4 Decoder wang2024exploring - - - - 59.2 - - - LPT deng2009imagenet - - - - 76.1 - - 79.3 Linear Probing 17.5 13.3 21.1 16.7 60.4 48.9 60.0 63.9 Full-FT 38.0 43.8 41.5 20.0 76.1 75.7 76.9 75.3 cRT kang2019decoupling 37.7 42.9 39.3 25.0 44.4 51.0 43.1 35.4 PEFT shi2024long 38.5 43.3 40.4 25.5 79.1 72.4 79.0 81.1 Model Soups wortsman2022model 38.0 45.6 40.2 20.0 76.4 77.1 76.8 75.6 LT-Soups (Ours) 39.3 42.4 40.7 30.9 78.2 76.7 78.5 78.2
@@ 5.3 Empirical analysis of LT-Soups
LT-Soups averages the weights of $NM$ models, where $N$ is the number of subsets used during each fine-tuning and $M$ is the number of bootstraps per subset. For TinyImageNet-LT (imbalance ratio 100 and head-tail ratio 0.3), we use NN=8 and MM=2. To show the impact of our proposed weight averaging scheme, we compare this with Soups-$\rho_{n}$ baselines that follow the same two-stage framework as LT-Soups, except in the first stage, they average 16 models, all trained on subsets with the same imbalance ratio $\rho_{n}$. Notably, Soups-100 aligns with the traditional model soup approach wortsman2022model, where the weights of 16 fully fine-tuned models on the entire dataset are averaged. As shown in Table 3, all of the soups baselines consistently outperform full fine-tuning, regardless of the subset choice. However, results show that different imbalance ratios yield varying outcomes across head and tail categories. For example, Soups-8 achieves the highest tail accuracy of 75.0, whereas Soups-100 reaches the highest head accuracy of 85.9. Rather than optimizing for a single imbalance ratio, LT-Soups applies weight averaging across the full spectrum, effectively merging the advantages of both approaches to achieve a more balanced overall trade-off. (See Table 10 in the Appendix for a similar analysis on PEFT.)
@@
By default, LT-Soups exhibits partial insensitivity to the choice of loss function, due to the balance introduced through subsampling and model averaging. However, in the first stage, each subset remains mildly imbalanced, though significantly less so than the full training set. Table 6 reports the impact of different class-balancing strategies used during this stage, including logit adjustment loss (LA), cross-entropy loss (CE), and class-balanced sampling (CB) kang2019decoupling. Unlike PEFT, which heavily depends on LA loss for optimal performance, LT-Soups is only moderately affected by the choice of class-balancing strategies, owing to the structural regularization introduced by training on multiple, complementary subsets.
@@ 6 Limitations and Future Work
While we define the head-tail ratio using a fixed sample threshold, this binary split may oversimplify the class distribution structure. A more nuanced approach could leverage the generalized Pareto distribution rootzen2006multivariate to model imbalance with controllable location, scale, and shape parameters. We leave exploration of such parameterized formulations for future work.
@@ Acknowledgements.
Broader Impact
Our work advances the field of long-tailed recognition by improving model performance across imbalanced datasets, which are prevalent in real-world applications such as medical imaging, wildlife monitoring, and autonomous driving. By enhancing accuracy for both head and tail classes, our method promotes fairness and inclusivity in AI systems, reducing biases toward dominant categories.
@@ Appendix A Additional ablations
An inherent advantage of model merging methods is their ability to improve prediction calibration metrics. We evaluate LT-Soups against PEFT and Full-FT by measuring Negative Log-Likelihood (NLL), Expected Calibration Error (ECE) naeini2015obtaining, and Brier score brier1950verification. For NLL and Brier scores, we also provide category-wise results. All metrics are computed after temperature tuning on the validation set. As shown in 7, LT-Soups consistently outperforms the other methods on TinyImageNet-LT in terms of calibration.
@@
A natural question is whether PEFT methods can replace the full fine-tuning process in LT-Soups. To investigate this, we use LoRA as a representative approach. In the first stage, we freeze the CLIP pre-trained weights and tune LoRA parameters using the same subsets as LT-Soups. The LoRA parameters are combined with the pre-trained weights before applying our merging schema. Finally, we retrain the classifier using the LA loss. The performance on TinyImagenet-LT dataset ($77.1$ vs $77.2$) matches that of end-to-end LoRA training. We hypothesize this outcome is due to a phenomenon observed in LLM literature shuttleworth2024lora, where LoRA introduces high-ranking singular vectors (intruder dimensions) that are absent in full fine-tuning. While these models achieve comparable task performance, they adapt less robustly to sequential tasks and diverge from the pre-training distribution.
@@ C.1 Imbalanced Classification
Class imbalance mitigation strategies broadly involve oversampling minority classes chawla2002smote, subsampling majority classes liu2008exploratory, or reweighting losses he2009learning. Generative approaches such as DiffuLT shao2024diffult train a diffusion model on a long-tailed distribution and then uses it to generate a balanced proxy dataset for training the final classifier. Subsampling risks losing majority-class discriminative patterns, oversampling may overfit minority classes zhou2020bbn, and reweighting struggles in overparameterized networks zhai2022understanding. Recent advances like balanced softmax ren2020balanced and its generalization, logit adjustment loss (LA) menon2020long address these issues by enforcing larger margins for tail classes, bridging data imbalance with geometric regularization.
@@
Decoupled learning frameworks address class imbalance through sequential training phases: representation learning via instance-balanced sampling followed by classifier refinement using class-balanced strategies kang2019decoupling, zhang2021distribution. This paradigm assumes model biases primarily reside in the classifier layer, positing that head-tail performance gaps can be resolved through post-hoc classifier calibration izmailov2022feature, yang2023change. However, shi2024long show empirically that this assumption becomes invalid when fine-tuning foundation models, neglecting a tailored strategy for representation learning, degrades both head and tail class performance due to catastrophic forgetting of pre-trained features mukhoti2023fine.
@@
Ensemble methods address data imbalance by combining specialized experts trained on complementary distributions cai2021ace, li2022nested. Notable approaches include: BBN’s dual-branch architecture balancing original and re-sampled distributions zhou2020bbn; RIDE’s dynamic routing of instances to distribution-aware experts wang2020long; and LFME’s multi-teacher distillation across many/medium/few-shot groups xiang2020learning; Reflective Learning zhao2024ltrl promotes consistency across training iterations by minimizing KL divergence between predictions and soft labels induced from feature similarity. While effective, these methods rely on heuristic expert specialization rules and often result in cumbersome architectures that hinder adaptation to foundation models, increase training complexity, and limit inference speed. Our work circumvents these limitations through two key innovations: (1) replacing specialized expert design with parallel fine-tuning of foundation models on controlled subsamples, and (2) employing model averaging and EMA instead of complex aggregation mechanisms. This preserves the ensemble’s variance-reduction benefits while maintaining the original foundation model’s architectural simplicity and computational efficiency.
@@ C.2 Model Merging
Model merging, also sometimes referred to to weight averaging, has gained significant attention in recent years as a promising research direction li2023deep, focusing on reducing communication costs in federated learning mcmahan2017communication and distributed training douillard2023diloco, enabling the efficient combination of multiple models without additional training ilharco2023editing, and enhancing model robustness in out-of-distribution scenarios wortsman2022model, rame2022diverse. Early approaches like Exponential Moving Average (EMA) tarvainen2017mean and Stochastic Weight Averaging (SWA) izmailov2018averaging have been widely adopted to accelerate training convergence, stability, and enhance the generalization capabilities of deep neural networks. Recent work extends merging to sequential adaptation: alexandrov-etal-2024-mitigating mitigates catastrophic forgetting in continual pretraining via iterative merging, while rame2024warp align LLMs through multi-stage averaging during RLHF. To our knowledge, no prior work applies model merging to imbalanced recognition. Unlike existing sequential merging approaches, our framework trains multiple models in parallel on complementary subsampled distributions, a critical design choice for handling long-tailed data. We propose the first schema specifically tailored for imbalance, integrating subsampling (to retain tail-class discriminability) and bootstrapping (to stabilize head-class representations). This parallelized merging strategy directly addresses feature-space asymmetry in long-tailed distributions while maintaining computational efficiency, enabling foundation models to adapt to extreme imbalance without sacrificing pre-trained generalization.
@@
We use CLIP with the ViT-B/16 backbone. Following radford2021learning, we adopt a prototypical classification head for $g$, where both features and classifier weights are $l_{2}$-normalized, and a temperature is applied to the logits. The parameters $\omega$ are initialized by generating text. We use descriptive prompts such as “a photo of a cat” or “a photo of a dog” to represent each class radford2021learning. for the classes and extracting corresponding textual features using the CLIP text encoder.
We optimize the model using the AdamW optimizer loshchilov2018decoupled. The batch size is set to 128, with learning rates of 3e43e-4 for both the representation and the classification stage. A cosine decay learning rate scheduler is employed, gradually reducing the learning rate to $0.1\cdot max\_lr$ after a warmup period spanning $max(100,0.01\cdot total\_steps)$ steps. The validation set of each dataset is used to select the best checkpoint. Table 11 shows the hyperparameters we used for each dataset. We select $N$ and $\lambda$ based on the validation set of each dataset and fix MM=2 across all experiments. We report all baseline results without test-time augmentation, which offers orthogonal gains.
@@
Dataset Classes Total samples Max samples Min samples $\rho$ $\eta$ CIFAR100-LT cao2019learning 100 10.8k 500 5 100 0.54 TinyImageNet-LT Le2015TinyIV 200 21.5k 500 5 100 0.53 Places-LT liu2019large 365 62.5k 4980 5 996 0.55 ImageNet-LT liu2019large 1000 115.8k 1280 5 256 0.62 iNaturalist van2018inaturalist 8,142 437.5k 1000 2 500 0.11 NIH-CXR-LT Holste2022LongTailedCO 20 88,637 53260 12 6491 5.66
@@ D.1 Full results
LDAM (Cao et al., 2019) ResNet-32 42.0 - - -
BBN (Zhou et al., 2020) ResNet-32 42.6 - - -
DiVE (He et al., 2021) ResNet-32 45.4 - - -
MiSLAS (Zhong et al., 2021) ResNet-32 47.0 - - -
BS (Ren et al., 2020) ResNet-32 50.8 - - -
PaCo (Cui et al., 2021) ResNet-32 52.0 - - -
BCL (Zhu et al., 2022) ResNet-32 51.9 - - -
@@
cRT kang2019decoupling ViT-B/16 78.8 89.7 79.7 65.1
PEFT shi2024long ViT-B/16 81.3 85.2 80.9 77.1
Model Soups wortsman2022model ViT-B/16 82.1 89.9 82.2 73.0
@@
OLTR liu2019large ResNet-152 35.9 44.7 37.0 25.3
cRT kang2019decoupling ResNet-152 36.7 42.0 37.6 26.4
LWS kang2019decoupling ResNet-152 37.6 40.6 39.1 28.6
MiSLAS (Zhong et al., 2021) ResNet-152 40.4 39.6 43.3 36.1
DisAlign (Zhang et al., 2021) ResNet-152 39.3 40.4 39.4 32.9
ALA (Zhao et al., 2022) ResNet-152 41.2 36.1 47.9 35.3
PaCo (Cui et al., 2021) ResNet-152 40.5 33.7 44.4 35.3
LiVT (Xu et al., 2023) ViT-B/16 40.8 48.1 40.6 27.5
@@
cRT kang2019decoupling ViT-B/16 44.4 51.0 43.1 35.4
BALLAD ma2021simple ViT-B/16 49.5 49.3 50.2 48.4
Decoder wang2024exploring ViT-B/16 46.8 - - -
LPT dong2022lpt ViT-B/16 50.1 49.3 52.3 46.9
@@
cRT kang2019decoupling ViT-B/16 44.4 51.0 43.1 35.4
PEFT shi2024long ViT-B/16 51.5 51.3 52.2 50.5
Model Soups wortsman2022model ViT-B/16 49.4 51.7 50.0 43.7
@@
cRT kang2019decoupling ResNet-50 47.3 58.8 44.0 26.1
LWS kang2019decoupling ResNet-50 47.7 57.1 45.2 29.3
MiSLAS (Zhong et al., 2021) ResNet-50 52.7 62.9 50.7 31.0
LA menon2020long ResNet-50 51.1 - - -
DisAlign (Zhang et al., 2021) ResNet-50 52.9 61.3 52.2 31.4
BCL (Zhu et al., 2022) ResNet-50 56.0 - - -
PaCo (Cui et al., 2021) ResNet-50 57.0 - - -
NCL (Li et al., 2022a) ResNet-50 57.4 - - -
LiVT (Xu et al., 2023) ViT-B/16 60.9 73.6 56.4 41.0
@@
BALLAD ma2021simple ViT-B/16 75.7 79.1 74.5 69.8
Decoder wang2024exploring ViT-B/16 73.2 - - -
@@
cRT kang2019decoupling ViT-B/16 72.6 81.1 70.6 56.1
PEFT shi2024long ViT-B/16 77.0 80.2 76.1 71.5
Model Soups wortsman2022model ViT-B/16 76.0 81.5 74.5 65.5
@@
cRT kang2019decoupling ResNet-50 38.0 43.3 37.4 30.0
LWS kang2019decoupling ResNet-50 28.0 45.7 23.0 08.3
CB LDAM-DRW cao2019learning ResNet-50 37.7 47.6 35.6 25.0
CB Softmax cui2019class ResNet-50 33.3 29.5 41.5 21.7
@@
cRT kang2019decoupling ViT-B/16 37.7 42.9 39.3 25.0
PEFT shi2024long ViT-B/16 38.5 43.3 40.4 25.5
Model Soups wortsman2022model ViT-B/16 38.0 45.6 40.2 20.0
@@
cRT kang2019decoupling ResNet-50 65.2 69.0 66.0 63.2
LWS kang2019decoupling ResNet-50 65.9 65.0 66.3 65.5
MiSLAS (Zhong et al., 2021) ResNet-50 71.6 73.2 72.4 70.4
DiVE (He et al., 2021) ResNet-50 69.1 70.6 70.0 67.7
DisAlign (Zhang et al., 2021) ResNet-50 69.5 69.1 69.9 69.4
ALA (Zhao et al., 2022) ResNet-50 69.6 69.5 70.2 69.0
RIDE (Wang et al., 2021c) ResNet-50 71.5 72.4 73.1 70.4
RIDE+CR (Ma et al., 2023) ResNet-50 73.5 74.0 74.3 73.1
RIDE+OTmix (Gao et al., 2023) ResNet-50 73.7 74.1 75.2 72.8
BCL (Zhu et al., 2022) ResNet-50 71.8 - - -
PaCo (Cui et al., 2021) ResNet-50 73.2 70.4 72.8 75.8
NCL (Li et al., 2022a) ResNet-50 74.2 72.0 74.9 73.8
GML (Suh & Seo, 2023) ResNet-50 74.5 - - -
LiVT (Xu et al., 2023) ViT-B/16 76.1 78.9 76.5 74.8
@@
PEFT shi2024long ViT-B/16 79.1 72.4 79.0 81.1
Model Soups wortsman2022model ViT-B/16 76.4 77.1 76.8 75.6

### v2

@@ 5 Experiments
We evaluate our method on both synthetically constructed and naturally occurring long-tailed (LT) datasets. For synthetic benchmarks, we use CIFAR-100-LT, ImageNet-LT, and Places-LT—long-tailed variants derived from their balanced counterparts by sampling class instances according to Pareto or exponential decay distributions [34]. These datasets exhibit sample counts ranging from 1,280 to as few as 5 images per class. For real-world evaluation, we include iNaturalist 2018 (8,142 classes, 437.5K images) and NIH-CXR-LT (20 classes, 88.5K images), which reflect different imbalance structures, with approximately 10% and 90% head classes, respectively. To assess performance across the long-tail spectrum, we also report the average accuracy across all five datasets. Following [34], we evaluate separately on many-shot (>>100 samples), medium-shot (20–100), and few-shot (<<20) class subsets. For ablation analysis, we use TinyImageNet-LT, which contains 200 classes with sample counts ranging from 500 in head classes to 5 in tail classes. To conserve space, we present only CLIP-based results in the main text; additional implementation details and extended results are included in Appendix D.
@@
Methods CIFAR100-LT Places-LT ImageNet-LT $\rho$=100 $\eta$=0.54 $\rho$=996 $\eta$=0.55 $\rho$=256 $\eta$=0.62 All Many Med. Few All Many Med. Few All Many Med. Few BALLAD [37] - - - - 49.5 49.3 50.2 48.4 75.7 79.1 74.5 69.8 Decoder [58] - - - - 46.8 - - - 73.2 - - - LPT [13] - - - - 50.1 49.3 52.3 46.9 - - - - Linear Probing 70.0 77.2 71.1 60.4 48.8 48.8 49.7 47.1 74.2 77.8 73.3 67.4 Full-FT 79.6 88.1 79.9 69.3 46.6 49.9 46.3 41.4 73.9 79.8 71.9 63.9 cRT [25] 78.8 89.7 79.7 65.1 44.4 51.0 43.1 35.4 72.6 81.1 70.6 56.1 PEFT [50] 81.3 85.2 80.9 77.1 51.5 51.3 52.2 50.5 77.0 80.2 76.1 71.5 Model Soups [60] 82.1 89.9 82.2 73.0 49.4 51.7 50.0 43.7 76.0 81.5 74.5 65.5 LT-Soups (Ours) 83.5 88.2 83.5 78.0 51.7 51.2 52.8 50.3 77.4 81.2 76.1 70.7
@@
Methods NIH-CXR-LT iNaturalist 2018 $\rho$=6491 $\eta$=5.66 $\rho$=500 $\eta$=0.11 All Many Med. Few All Many Med. Few BALLAD [37] 34.5 36.7 38.9 20.8 49.5 49.3 50.2 48.4 Decoder [58] - - - - 59.2 - - - LPT [12] - - - - 76.1 - - 79.3 Linear Probing 17.5 13.3 21.1 16.7 60.4 48.9 60.0 63.9 Full-FT 38.0 43.8 41.5 20.0 76.1 75.7 76.9 75.3 cRT [25] 37.7 42.9 39.3 25.0 44.4 51.0 43.1 35.4 PEFT [50] 38.5 43.3 40.4 25.5 79.1 72.4 79.0 81.1 Model Soups [60] 38.0 45.6 40.2 20.0 76.4 77.1 76.8 75.6 LT-Soups (Ours) 39.3 42.4 40.7 30.9 78.2 76.7 78.5 78.2
@@ 5.3 Empirical analysis of LT-Soups
LT-Soups averages the weights of $NM$ models, where $N$ is the number of subsets used during each fine-tuning and $M$ is the number of bootstraps per subset. For TinyImageNet-LT (imbalance ratio 100 and head-tail ratio 0.3), we use NN=8 and MM=2. To show the impact of our proposed weight averaging scheme, we compare this with Soups-$\rho_{n}$ baselines that follow the same two-stage framework as LT-Soups, except in the first stage, they average 16 models, all trained on subsets with the same imbalance ratio $\rho_{n}$. Notably, Soups-100 aligns with the traditional model soup approach [60], where the weights of 16 fully fine-tuned models on the entire dataset are averaged. As shown in Table 3, all of the soups baselines consistently outperform full fine-tuning, regardless of the subset choice. However, results show that different imbalance ratios yield varying outcomes across head and tail categories. For example, Soups-8 achieves the highest tail accuracy of 75.0, whereas Soups-100 reaches the highest head accuracy of 85.9. Rather than optimizing for a single imbalance ratio, LT-Soups applies weight averaging across the full spectrum, effectively merging the advantages of both approaches to achieve a more balanced overall trade-off. (See Table 10 in the Appendix for a similar analysis on PEFT.)
@@
By default, LT-Soups exhibits partial insensitivity to the choice of loss function, due to the balance introduced through subsampling and model averaging. However, in the first stage, each subset remains mildly imbalanced, though significantly less so than the full training set. Table 6 reports the impact of different class-balancing strategies used during this stage, including logit adjustment loss (LA), cross-entropy loss (CE), and class-balanced sampling (CB) [25]. Unlike PEFT, which heavily depends on LA loss for optimal performance, LT-Soups is only moderately affected by the choice of class-balancing strategies, owing to the structural regularization introduced by training on multiple, complementary subsets.
@@ 6 Limitations and Future Work
While we define the head-tail ratio using a fixed sample threshold, this binary split may oversimplify the class distribution structure. A more nuanced approach could leverage the generalized Pareto distribution [47] to model imbalance with controllable location, scale, and shape parameters. We leave exploration of such parameterized formulations for future work.
@@ Appendix A Additional ablations
An inherent advantage of model merging methods is their ability to improve prediction calibration metrics. We evaluate LT-Soups against PEFT and Full-FT by measuring Negative Log-Likelihood (NLL), Expected Calibration Error (ECE) [42], and Brier score [3]. For NLL and Brier scores, we also provide category-wise results. All metrics are computed after temperature tuning on the validation set. As shown in 7, LT-Soups consistently outperforms the other methods on TinyImageNet-LT in terms of calibration.
@@
A natural question is whether PEFT methods can replace the full fine-tuning process in LT-Soups. To investigate this, we use LoRA as a representative approach. In the first stage, we freeze the CLIP pre-trained weights and tune LoRA parameters using the same subsets as LT-Soups. The LoRA parameters are combined with the pre-trained weights before applying our merging schema. Finally, we retrain the classifier using the LA loss. The performance on TinyImagenet-LT dataset ($77.1$ vs $77.2$) matches that of end-to-end LoRA training. We hypothesize this outcome is due to a phenomenon observed in LLM literature [51], where LoRA introduces high-ranking singular vectors (intruder dimensions) that are absent in full fine-tuning. While these models achieve comparable task performance, they adapt less robustly to sequential tasks and diverge from the pre-training distribution.
@@ C.1 Imbalanced Classification
Class imbalance mitigation strategies broadly involve oversampling minority classes [7], subsampling majority classes [33], or reweighting losses [17]. Generative approaches such as DiffuLT [48] train a diffusion model on a long-tailed distribution and then uses it to generate a balanced proxy dataset for training the final classifier. Subsampling risks losing majority-class discriminative patterns, oversampling may overfit minority classes [74], and reweighting struggles in overparameterized networks [66]. Recent advances like balanced softmax [46] and its generalization, logit adjustment loss (LA) [40] address these issues by enforcing larger margins for tail classes, bridging data imbalance with geometric regularization.
@@
Decoupled learning frameworks address class imbalance through sequential training phases: representation learning via instance-balanced sampling followed by classifier refinement using class-balanced strategies [25, 67]. This paradigm assumes model biases primarily reside in the classifier layer, positing that head-tail performance gaps can be resolved through post-hoc classifier calibration [23, 64]. However, [50] show empirically that this assumption becomes invalid when fine-tuning foundation models, neglecting a tailored strategy for representation learning, degrades both head and tail class performance due to catastrophic forgetting of pre-trained features [41].
@@
Ensemble methods address data imbalance by combining specialized experts trained on complementary distributions [4, 31]. Notable approaches include: BBN’s dual-branch architecture balancing original and re-sampled distributions [74]; RIDE’s dynamic routing of instances to distribution-aware experts [57]; and LFME’s multi-teacher distillation across many/medium/few-shot groups [62]; Reflective Learning [69] promotes consistency across training iterations by minimizing KL divergence between predictions and soft labels induced from feature similarity. While effective, these methods rely on heuristic expert specialization rules and often result in cumbersome architectures that hinder adaptation to foundation models, increase training complexity, and limit inference speed. Our work circumvents these limitations through two key innovations: (1) replacing specialized expert design with parallel fine-tuning of foundation models on controlled subsamples, and (2) employing model averaging and EMA instead of complex aggregation mechanisms. This preserves the ensemble’s variance-reduction benefits while maintaining the original foundation model’s architectural simplicity and computational efficiency.
@@ C.2 Model Merging
Model merging, also sometimes referred to to weight averaging, has gained significant attention in recent years as a promising research direction [32], focusing on reducing communication costs in federated learning [39] and distributed training [15], enabling the efficient combination of multiple models without additional training [22], and enhancing model robustness in out-of-distribution scenarios [60, 45]. Early approaches like Exponential Moving Average (EMA) [54] and Stochastic Weight Averaging (SWA) [24] have been widely adopted to accelerate training convergence, stability, and enhance the generalization capabilities of deep neural networks. Recent work extends merging to sequential adaptation: Alexandrov et al. [1] mitigates catastrophic forgetting in continual pretraining via iterative merging, while Ramé et al. [44] align LLMs through multi-stage averaging during RLHF. To our knowledge, no prior work applies model merging to imbalanced recognition. Unlike existing sequential merging approaches, our framework trains multiple models in parallel on complementary subsampled distributions, a critical design choice for handling long-tailed data. We propose the first schema specifically tailored for imbalance, integrating subsampling (to retain tail-class discriminability) and bootstrapping (to stabilize head-class representations). This parallelized merging strategy directly addresses feature-space asymmetry in long-tailed distributions while maintaining computational efficiency, enabling foundation models to adapt to extreme imbalance without sacrificing pre-trained generalization.
@@
We use CLIP with the ViT-B/16 backbone. Following [43], we adopt a prototypical classification head for $g$, where both features and classifier weights are $l_{2}$-normalized, and a temperature is applied to the logits. The parameters $\omega$ are initialized by generating text. We use descriptive prompts such as “a photo of a cat” or “a photo of a dog” to represent each class [43]. for the classes and extracting corresponding textual features using the CLIP text encoder.
We optimize the model using the AdamW optimizer [36]. The batch size is set to 128, with learning rates of 3e43e-4 for both the representation and the classification stage. A cosine decay learning rate scheduler is employed, gradually reducing the learning rate to $0.1\cdot max\_lr$ after a warmup period spanning $max(100,0.01\cdot total\_steps)$ steps. The validation set of each dataset is used to select the best checkpoint. Table 11 shows the hyperparameters we used for each dataset. We select $N$ and $\lambda$ based on the validation set of each dataset and fix MM=2 across all experiments. We report all baseline results without test-time augmentation, which offers orthogonal gains.
@@
Dataset Classes Total samples Max samples Min samples $\rho$ $\eta$ CIFAR100-LT [5] 100 10.8k 500 5 100 0.54 TinyImageNet-LT [30] 200 21.5k 500 5 100 0.53 Places-LT [34] 365 62.5k 4980 5 996 0.55 ImageNet-LT [34] 1000 115.8k 1280 5 256 0.62 iNaturalist [56] 8,142 437.5k 1000 2 500 0.11 NIH-CXR-LT [19] 20 88,637 53260 12 6491 5.66
@@ D.1 Full results
LDAM [5] ResNet-32 42.0 - - -
BBN [74] ResNet-32 42.6 - - -
DiVE [18] ResNet-32 45.4 - - -
MiSLAS [72] ResNet-32 47.0 - - -
BS [46] ResNet-32 50.8 - - -
PaCo [10] ResNet-32 52.0 - - -
BCL [75] ResNet-32 51.9 - - -
@@
cRT [25] ViT-B/16 78.8 89.7 79.7 65.1
PEFT [50] ViT-B/16 81.3 85.2 80.9 77.1
Model Soups [60] ViT-B/16 82.1 89.9 82.2 73.0
@@
OLTR [34] ResNet-152 35.9 44.7 37.0 25.3
cRT [25] ResNet-152 36.7 42.0 37.6 26.4
LWS [25] ResNet-152 37.6 40.6 39.1 28.6
MiSLAS [72] ResNet-152 40.4 39.6 43.3 36.1
DisAlign [67] ResNet-152 39.3 40.4 39.4 32.9
ALA [71] ResNet-152 41.2 36.1 47.9 35.3
PaCo [10] ResNet-152 40.5 33.7 44.4 35.3
LiVT [63] ViT-B/16 40.8 48.1 40.6 27.5
@@
cRT [25] ViT-B/16 44.4 51.0 43.1 35.4
BALLAD [37] ViT-B/16 49.5 49.3 50.2 48.4
Decoder [58] ViT-B/16 46.8 - - -
LPT [13] ViT-B/16 50.1 49.3 52.3 46.9
@@
cRT [25] ViT-B/16 44.4 51.0 43.1 35.4
LIFT [50] ViT-B/16 51.5 51.3 52.2 50.5
Model Soups [60] ViT-B/16 49.4 51.7 50.0 43.7
@@
cRT [25] ResNet-50 47.3 58.8 44.0 26.1
LWS [25] ResNet-50 47.7 57.1 45.2 29.3
MiSLAS [72] ResNet-50 52.7 62.9 50.7 31.0
LA [40] ResNet-50 51.1 - - -
DisAlign [67] ResNet-50 52.9 61.3 52.2 31.4
BCL [75] ResNet-50 56.0 - - -
PaCo [10] ResNet-50 57.0 - - -
NCL [31] ResNet-50 57.4 - - -
LiVT [63] ViT-B/16 60.9 73.6 56.4 41.0
@@
BALLAD [37] ViT-B/16 75.7 79.1 74.5 69.8
Decoder [58] ViT-B/16 73.2 - - -
@@
cRT [25] ViT-B/16 72.6 81.1 70.6 56.1
LIFT [50] ViT-B/16 77.0 80.2 76.1 71.5
Model Soups [60] ViT-B/16 76.0 81.5 74.5 65.5
@@
cRT [25] ResNet-50 38.0 43.3 37.4 30.0
LWS [25] ResNet-50 28.0 45.7 23.0 08.3
CB LDAM-DRW [5] ResNet-50 37.7 47.6 35.6 25.0
CB Softmax [11] ResNet-50 33.3 29.5 41.5 21.7
@@
cRT [25] ViT-B/16 37.7 42.9 39.3 25.0
LIFT [50] ViT-B/16 38.5 43.3 40.4 25.5
Model Soups [60] ViT-B/16 38.0 45.6 40.2 20.0
@@
cRT [25] ResNet-50 65.2 69.0 66.0 63.2
LWS [25] ResNet-50 65.9 65.0 66.3 65.5
MiSLAS [72] ResNet-50 71.6 73.2 72.4 70.4
DiVE [18] ResNet-50 69.1 70.6 70.0 67.7
DisAlign [67] ResNet-50 69.5 69.1 69.9 69.4
ALA [71] ResNet-50 69.6 69.5 70.2 69.0
RIDE [57] ResNet-50 71.5 72.4 73.1 70.4
RIDE+CR [38] ResNet-50 73.5 74.0 74.3 73.1
RIDE+OTmix [16] ResNet-50 73.7 74.1 75.2 72.8
BCL [75] ResNet-50 71.8 - - -
PaCo [10] ResNet-50 73.2 70.4 72.8 75.8
NCL [31] ResNet-50 74.2 72.0 74.9 73.8
GML [52] ResNet-50 74.5 - - -
LiVT [63] ViT-B/16 76.1 78.9 76.5 74.8
@@
LIFT [50] ViT-B/16 79.1 72.4 79.0 81.1
Model Soups [60] ViT-B/16 76.4 77.1 76.8 75.6